OpenStudy (anonymous):

Would you make more interest by investing five thousand dollars in a 6-month certificate of deposit (CD) at six percent interest or by investing three thousand dollars in a twelve-month CD at four percent interest? Show your work.

OpenStudy (anonymous):

If the interest rates are per annum then then the basic equation would be A=P(1+r)t which for the first case would be A=5000(1.06)0.5 and the Second case B=3000(1.04)1 A = $5,147.81 B= $3,120.00